# Isotopy classification of Morse polynomials of degree three on ${\mathbb{R}}^{3}$

### v8

@@ 3. For any $t>0$, the action
All eight virtual components mentioned in Proposition 40 are achiral. In particular, there are exactly eight isotopy classes of degree three Morse polynomials ${\mathbb{R}}^{3}\to{\mathbb{R}}$ with non-discriminantal principal homogeneous parts of type $\Xi_{2}$ and eight real critical points.
Proof. Consider the virtual component, whose D-graph is shown in Fig. 11. By Lemma 22, there is a virtual Morse function in it, whose critical values are ordered in ${\mathbb{R}}^{1}$ as follows: $\begin{array}[]{cccc}4&5&7&8\\ 1&2&3&6\end{array}$ (in this picture we refer to the positions of the corresponding vertices of the D-graph of Fig. 11). Let $f$ be an arbitrary generic polynomial of the form (7) with only positive critical values, associated with this virtual Morse function. Vertices 2 and 3 are not connected by edges of our D-graph, nor are vertices 7 and 8. Therefore, there exists a path in the real subspace of the parameter space $\Theta$ of the deformation (7), starting from $f$, along which only two surgeries of type s2 occur, permuting the critical values of the corresponding pairs of critical points, and all the remaining four critical values are fixed along this path. The virtual Morse function of the resulting generic polynomial $\tilde{f}$ coincides with that of the initial polynomial $f$, so by Theorem 12 either $f$ and $\tilde{f}$ are isotopic to each other in the space of generic polynomials ${\mathbb{R}}^{3}\to{\mathbb{R}}$, or their isotopy classes are mapped to each other by the reflection of arguments in an arbitrary plane in ${\mathbb{R}}^{3}$.
However, these polynomials cannot be isotopic in this space. Indeed, each isotopy class of generic polynomials with four critical points of Morse index 1 defines an orientation of ${\mathbb{R}}^{3}$ in the following way. We number these points in ascending order of their critical values. By Lemma 24 these points are the vertices of a simplex in ${\mathbb{R}}^{3}$, and their order defines an orientation of this simplex and of the whole space. This orientation is constant along the isotopy classes of generic polynomials. Given a generic path in an arbitrary isotopy class of Morse polynomials, these canonical orientations of the generic polynomials at the beginning and at the end of this path are the same if and only if the number of intersection points of this path with the set of non-generic Morse polynomials corresponding to surgeries of type s2, which transpose critical values of two critical points with Morse index 1, is even. But the path constructed above provides only one such transposition.
@@ 5.2. Polynomials with six critical points
Let us realize these components. The critical point of class $D_{5}$ of the polynomial (41) with $\lambda>0$ can be decomposed by an arbitrarily small perturbation into three Morse critical points of Morse index 11 and either two critical points of index $2$ or two points of indices 0 and 22 (see two parts of Fig. 15). If we also split the point of class $A_{3}$ of the same polynomial into one real Morse critical point (which necessarily has Morse index 2) and two non-real points, we obtain (depending on the previous choice) polynomials with passports $(0,3,3,0)$ and $(1,3,2,0)$. For example, these perturbations of the polynomial (41) with $\lambda=\frac{1}{4}$ are realized by polynomials
@@
Two polynomials with passport $(0,3,3,0)$ constructed in the previous two paragraphs have different values of the invariant Card, in particular they belong ti different isotopy classes of Morse polynomials.
Proof. These polynomials are obtained by single Morse surgeries of type s1 from polynomials with eight real critical points, whose D-graphs are shown in Figs. 11 and 14, respectively. The virtual Morse functions associated with these polynomials are derived by the standard formulas (implemented in our program) from the similar data of the initial polynomials. These virtual Morse functions are
@@ 5.4. Polynomials with two critical points
(52) $x^{3}-xz^{2}+y^{2}z+\lambda z^{2}+\varepsilon x,\ \ \varepsilon>0$
(which is also invariant under the reflection in the plane $\{y=0\}$).
@@ 6. Isotopy classification of strictly Morse polynomials with eight real critical points
Corollary 52.
There are exactly 842 ((respectively, 1416)) isotopy classes of strictly Morse polynomials of type $\Xi_{1}$ ((respectively, $\Xi_{2})$ with eight real critical points. $\Box$
Proof of Theorem 51. For any generic polynomial $f$, denote by $f^{\uparrow}$ an arbitrary polynomial of the form $f+c$, where $c$ is a constant such that all real critical values of $f+c$ are positive.
@@
Consider an arbitrary virtual component $\Omega$, all elements of which have eight real critical points. By statement 1 of Lemma 53, the number of isotopy classes of strictly Morse polynomials with set-valued invariant $\Omega$ is equal to the number of isotopy classes of such polynomials with only positive critical values.
@@ 7.1. Normalization of D-graphs
Consider one of the D-graphs of Figs. 59 and 1114, whose vertices are marked by integer Morse indices of the corresponding critical points (determined in Propositions 31 and 41).
@@
An edge of a D-graph with vertices marked by integer Morse indices is called normal if it is directed from a vertex with a smaller Morse index to a vertex with a larger one, and is called tunnel otherwise. The normalization of a D-graph consists in removing all its tunnel edges.
@@
A very similar situation arises for D-graphs of degree four polynomials ${\mathbb{R}}^{2}\to{\mathbb{R}}$, see [31]; in particular, the extended Coxeter-Dynkin graph of type $\tilde{E}_{7}$ appears there.
@@ 7.2. D-graphs and Morse complexes
over ${\mathbb{Z}}_{2}$, where the generators of $C_{i}$ correspond to the vertices marked by $i$, and the matrix coefficient connecting two such generators of neighboring groups $C_{i}$ is equal to the parity of the number of segments drawn between these vertices in the D-graph.
@@ 8. Proof of Theorems 2 and 3
The $E_{7}$ singularities do not appear in the decompositions of the $P_{8}$ singularities (even in the complex domain), which excludes the multisingularity $E_{7}+A_{1}$ for both cases $\Xi_{1}$ and $\Xi_{2}$.
@@
Furthermore, none of the D-graphs shown in Figs. 1114 can be split into two ordered subgraphs, whose inner edges form canonical Coxeter-Dynkin graphs of class $D_{4}$, in such a way that all edges of the initial D-graph connecting the vertices of different subgraphs are directed from the first subgraph to the second. This prohibits the polynomials of class $\Xi_{2}$ with two critical points of class $D_{4}^{-}$. The absence of classes $D_{4}^{-}+D_{4}^{+}$, $D_{4}^{-}+A_{4}$, $D_{6}^{-}+A_{2}$ and $E_{7}+A_{1}$ for polynomials of any type $\Xi_{1}$ or $\Xi_{2}$ is already proved in the proof of Theorem 2. $\Box$

### v9

@@ 5.1. Polynomials with eight critical points
All eight virtual components mentioned in Proposition 43 are achiral. In particular, there are exactly eight isotopy classes of degree three Morse polynomials ${\mathbb{R}}^{3}\to{\mathbb{R}}$ with non-discriminant principal homogeneous parts of type $\Xi_{2}$ and eight real critical points.
Proof. Consider the virtual component, whose D-graph is shown in Fig. 11. By Lemma 25, there is a virtual Morse function in it, whose critical values are ordered in ${\mathbb{R}}^{1}$ as follows: $\begin{array}[]{cccc}4&8&5&7\\ 1&2&6&3\end{array}$ (in this picture we refer to the positions of the corresponding vertices of the D-graph of Fig. 11). Let $f$ be an arbitrary generic polynomial of the form (7) with only positive critical values, associated with this virtual Morse function. Vertices 2 and 3 are not connected by edges of our D-graph, nor are vertices 7 and 8. Therefore, there exists a path in the real subspace of the parameter space $\Theta$ of the deformation (7), starting from $f$, along which only two surgeries of type s2 occur, permuting the critical values of the corresponding pairs of critical points, and all the remaining four critical values are fixed along this path. The virtual Morse function of the resulting generic polynomial $\tilde{f}$ coincides with that of the initial polynomial $f$, so by Theorem 15 either $f$ and $\tilde{f}$ are isotopic to each other in the space of generic polynomials ${\mathbb{R}}^{3}\to{\mathbb{R}}$, or their isotopy classes are mapped to each other by the reflection of arguments in an arbitrary plane in ${\mathbb{R}}^{3}$.
However, these polynomials cannot be isotopic in this space. Indeed, each isotopy class of generic polynomials with four critical points of Morse index 1 defines an orientation of ${\mathbb{R}}^{3}$ in the following way. We number these points in ascending order of their critical values. By Lemma 27 these points are the vertices of a simplex in ${\mathbb{R}}^{3}$, and their order defines an orientation of this simplex and of the whole space. This orientation is constant along the isotopy classes of generic polynomials. Given a generic path in an arbitrary isotopy class of Morse polynomials, these canonical orientations of the generic polynomials at the beginning and at the end of this path are the same if and only if the number of intersection points of this path with the set of non-strictly Morse polynomials corresponding to surgeries of type s2, which transpose critical values of two critical points with Morse index 1, is even. But the path constructed above provides only one such transposition.
@@ 5.2. Polynomials with six critical points
Let us realize these components. The critical point of class $D_{5}$ of the polynomial (41) with $\lambda>0$ can be decomposed by an arbitrarily small perturbation into three Morse critical points with Morse index 11 and either two critical points with index $2$ or two points with indices 0 and $2$ (see two parts of Fig. 15). If we also split the point of class $A_{3}$ of the same polynomial into one real Morse critical point (which necessarily has Morse index 2) and two non-real points, then we obtain (depending on the previous choice) polynomials with passports $(0,3,3,0)$ and $(1,3,2,0)$. For example, these perturbations of the polynomial (41) with $\lambda=\frac{1}{4}$ are realized by polynomials
@@
Two polynomials with passport $(0,3,3,0)$ constructed in the previous two paragraphs have different values of the invariant Card, in particular they belong to different isotopy classes of Morse polynomials.
Proof. These polynomials are obtained by single Morse surgeries of type s1 from polynomials with eight real critical points, whose D-graphs are shown in Figs. 11 and 14, respectively. The virtual Morse functions associated with these polynomials are derived using the standard formulas (implemented in our program) from the similar data of the initial polynomials. These virtual Morse functions are
@@ 5.4. Polynomials with two critical points
(52) $x^{3}-xz^{2}+y^{2}z+\lambda z^{2}+\varepsilon x,\ \ \varepsilon>0.$
This polynomial is also invariant under the reflection in the plane $\{y=0\}$, thus the virtual component is achiral and is associated with only one isotopy class of Morse polynomials.
@@ 6. Isotopy classification of strictly Morse polynomials with eight real critical points
The first statement of Theorem 2 is a direct corollary of this theorem and Proposition 17.
For any generic polynomial $f$, denote by $f^{\uparrow}$ an arbitrary polynomial of the form $f+c$, where $c$ is a constant such that all real critical values of $f+c$ are positive.
@@
Proof of Theorem 54. Consider an arbitrary virtual component $\Omega$, all elements of which have eight real critical points. By statement 1 of Lemma 55, the number of isotopy classes of strictly Morse polynomials with set-valued invariant $\Omega$ is equal to the number of isotopy classes of such polynomials with only positive critical values.
@@
Proof of the second statement of Theorem 2. Denote by $\mathfrak{G}$} the connected group of orientation preserving affine transformations $G:{\mathbb{R}}^{3}\to{\mathbb{R}}^{3}$. This group acts on the spaces $\Xi_{1}$ and $\Xi_{2}$ by mapping each polynomial $f$ to $f\circ G$. The action of $\mathfrak{G}$} on the set of strictly Morse polynomials with eight critical points is free. Indeed, if $f$ is such a polynomial and $f=f\circ G$, $G\in{\mathfrak{G}}$, then the transformation $G$ maps each critical point of $f$ to itself. These points affinely generate ${\mathbb{R}}^{3}$, so $G=\mbox{Id}$. According to the versality property of the deformation (7), the orbits of this action intersect the space of polynomials (7) transversally. Therefore, each connected component $U$ of the set of strictly Morse polynomials with eight critical points is a fiber bundle with fiber $\mathfrak{G}$}, whose base is an arbitrary connected component of the intersection of the component $U$ and the space (7). This intersection is diffeomorphic to ${\mathbb{R}}^{8}$. Indeed, the Lyashko–Looijenga map makes it a covering over the domain $\{x_{1}<x_{2}<\cdots<x_{8}\}\subset{\mathbb{R}}^{8}$. Therefore, component $U$ is a fiber bundle with a contractible base, and its fiber is homotopy equivalent to $SL(3,{\mathbb{R}})$. $\Box$
@@ 7.1. Normalization of D-graphs
Consider one of the D-graphs of Figs. 59 and 1114, whose vertices are marked by integer Morse indices of the corresponding critical points (determined in Propositions 34 and 44).
@@
An edge of a D-graph with vertices marked by integer Morse indices is called normal if
a) it is oriented from a vertex with a smaller Morse index to a vertex with a larger index,
b) it is dashed if the parities of Morse indices of the critical points corresponding to its ends are the same; it is solid if these parities are different.
Otherwise, this edge is called a tunnel edge. The normalization of a $D$-graph consists in removing all its tunnel edges.
@@
A very similar situation arises for D-graphs of degree four polynomials ${\mathbb{R}}^{2}\to{\mathbb{R}}$ and for perturbations of $J_{10}$ singularities, see [36] and [37]. In particular, the extended Coxeter-Dynkin graphs of types $\tilde{E}_{7}$ and $\tilde{E}_{8}$ appear there.
@@ 7.2. D-graphs and Morse complexes
over ${\mathbb{Z}}_{2}$, where the generators of $C_{i}$ correspond to the vertices marked by $i$, and the matrix coefficient connecting two such generators of neighboring groups $C_{i}$ and $C_{i-1}$ is equal to the parity of the number of segments drawn between these vertices in the D-graph.
@@ 8. Proof of Theorems 4 and 5
The $A_{7}$, $D_{7}$ and $E_{7}$ singularities do not appear in the decompositions of the $P_{8}$ singularities (even in the complex domain), which excludes the multisingularities $A_{7}+A_{1}$, $D_{7}+A_{1}$, and $E_{7}+A_{1}$ for both cases $\Xi_{1}$ and $\Xi_{2}$.
@@
Furthermore, none of the D-graphs shown in Figs. 1114 can be split into two ordered subgraphs, whose inner edges form canonical Coxeter-Dynkin graphs of class $D_{4}$, in such a way that all edges of the initial D-graph connecting the vertices of different subgraphs are directed from the first subgraph to the second. This prohibits the polynomials of class $\Xi_{2}$ with two critical points of class $D_{4}^{-}$. The absence of classes $D_{4}^{-}+D_{4}^{+}$, $D_{4}^{-}+A_{4}$, $D_{6}^{-}+A_{2}$, $A_{7}+A_{1}$, $D_{7}+A_{1}$, and $E_{7}+A_{1}$ for polynomials of any type $\Xi_{1}$ or $\Xi_{2}$ is already proved in the proof of Theorem 4. $\Box$
9. A few problems
1. List all the isotopy classes of Morse polynomials within the parameter space of the canonical miniversal deformation (7) of $P_{8}$ singularities. These classes are the connected components of the intersections of this space with the isotopy classes computed above. In general, there can be more than one such component. For example, in the case of the isotopy class of type $\Xi_{1}$ with D-graph shown in Fig. 8, the boundary of each such component contains only one polynomial of the form (7) with critical points of types $E_{6}$ and $A_{2}$ and corresponding critical values 0 and 1. According to Lemma 38, there are three such polynomials. One has the form (28), and the other two are obtained from it by coordinate permutation. Therefore, there are three different such components whose boundaries contain these three polynomials.
2. Study the topology of the isotopy classes of Morse polynomials. The homotopy types of the caustic complements of simple singularities of classes $A_{k}$, $D_{k}$, and $E_{6}$ were computed in [27], [28]. All of them are either contractible or homotopy equivalent to the circles. In [35], it was proven that the caustic complements of parabolic singularities of classes $P_{8}^{2}$ and $J_{10}^{3}$ have nontrivial second homology groups.
The following are some sources of homology classes of isotopy classes.
A. For each isotopy class with passport $(m_{0},m_{1},m_{2},m_{3})$, its map to the configuration space of collections of $m_{0}+m_{1}+m_{2}+m_{3}$ distinct points in ${\mathbb{R}}^{3}$ colored into four colors (or fewer if some of the $m_{i}$ are zero) is defined in the obvious way. Some cohomology classes can be induced from the cohomology group of this configuration space.
B. In the case of isotopy classes with non-real critical points, a ${\mathbb{Z}}_{2}$-valued 1-cohomology class is defined by the number of $s4$ surgeries occurring along the loops inside these classes. Can this cocycle be extended to an integer one? Can this cocycle take non-zero values on the isotopy classes of strictly Morse polynomials?
C. Two ${\mathbb{Z}}_{2}$-valued 1-cohomology classes are also defined by the numbers of $s2$ surgeries at which two critical values at the Morse points with the same Morse index (equal to 1 or 2) become equal.
D. If an isotopy class of Morse polynomials is achiral and does not contain polynomials that are invariant under the reflections in planes of ${\mathbb{R}}^{3}$, then it is the base of a nontrivial two-fold covering. Therefore, it admits a nontrivial 1-cohomology class with ${\mathbb{Z}}_{2}$ coefficients.
3. Study the subdivision of isotopy classes of Morse polynomials with non-real critical values by the complex Maxwell set, which consists of the polynomials having non-real critical points with real critical values.
4. The first version of our program listing the virtual functions of singularities was written in Fortran in 1984; see [6]. Revising and rewriting it it in a more advanced language would be useful.
The bottleneck of its computational complexity lies in the procedure of comparing new virtual functions with previously found ones. Clustering the formal graph by combining at one vertex all virtual functions that differ only in the number of negative critical values is likely to reduce computation time for a singularity with Milnor number $\mu$ by about a factor of $\mu^{2}$.